Human blood typically contains 1.04 kg/L of platelets. A 1.01 pints of blood would contain what mass (in grams) of platelets? (1 gallon = 3.785 L, 1 gallon = 8 pints).




1
Expert's answer
2021-09-09T02:36:02-0400

1 gallon = 8 pints

? 1.01

= 1.01/8 = 0.12625 gallon

1 gallon = 3.785 L

0.12625 = ?

= 3.785 × 0.12625 = 0.478 L

1L = 1000g

= 0.478 × 1000 = 478grams